OpenStudy (marshmallow_swag14):

What is the LCD of 5/6 and -3/8 ?

OpenStudy (marshmallow_swag14):

@lizzy_oh.hi can you help me? Please?

OpenStudy (mathstudent55):

What is the smallest number that you can divide evenly by 6 and 8?

OpenStudy (marshmallow_swag14):

2 right?

OpenStudy (mathstudent55):

No. That is the GCF, the greatest common factor. That is the largest number that divides evenly into 6 and 8.

OpenStudy (mathstudent55):

You want a number that is greater than or equal to 6 and greater than or equal to 8. When you divide that number by 6 or by 8 you get no remainder.

OpenStudy (mathstudent55):

For example, the LCD of 3 and 4 is 12 because 12/3 = 4 (no remainder) and 12/4 = 3 (no remainder).

OpenStudy (marshmallow_swag14):

oh ok

OpenStudy (mathstudent55):

You can list the multiples of both 6 and 8 until you find the first one they both have in common.

OpenStudy (marshmallow_swag14):

but it is negative...

OpenStudy (marshmallow_swag14):

or will that even matter?

OpenStudy (mathstudent55):

Write a list of 6 multiplied by 1, 2, 3, 4, 5 Write a list of 8 multiplied by 1, 2, 3, 4, 5 What is the smallest number in both lists of multiples?

OpenStudy (mathstudent55):

The number is positive.

OpenStudy (marshmallow_swag14):

ok thx and 40 is the lowest

OpenStudy (mathstudent55):

40 is not a multiple of 6, so 40 can't be the LCD. Here is a list of the first several multiples of 6 greater than 0: 6, 12, 18, 24, 30, 26, 42 Ok? It's simply 6 * 1, 6 * 2, 6 * 3, ..., 6 * 7 Now do the same for 8. What are the first few multiples of 8?

OpenStudy (marshmallow_swag14):

Oh HAHA I was thinking 5

OpenStudy (marshmallow_swag14):

its 24

OpenStudy (mathstudent55):

Correct. Multiples of 6: 6., 12, 18, 24, 30, 36, 42 Multiples of 8: 8, 16, 24, 32, 40 The smallest multiple 6 and 8 have in common is 24.

OpenStudy (marshmallow_swag14):

Thank you for the Help!!! :)

OpenStudy (mathstudent55):

You're welcome.